the sum &product of three numbers is same then what are those numbers?

Grade:12

5 Answers

kalyan
8 years ago
its simple yaaar XYZ = X+Y+Z
the only possibility is 1 ,2,3
sri tanish
8 years ago
the only possible numbers satisfying the given property are 1,2,3 .
VERIFICATION – 1+2+3=6
     1*2*3=6
pa1
8 years ago
1.2.3
SUM IS 1+2+3=6
PRODUCT IS 1×2×3=6
SHANMUKESHWAR
8 years ago
1,2,3
Sanny Chauhan
8 years ago
What about these number
 0, a, -a, ..
Where a \subset R
For example numbers are (0, 4, -4) , (0, 6, -6)
 
This is for completing 100 characters.